### v1

# SynCo: Synthetic Hard Negatives in Contrastive Learning for Better Unsupervised Visual Representations

Abstract
Contrastive learning has become a dominant approach in self-supervised visual representation learning, with hard negatives—samples that closely resemble the anchor—being key to enhancing the discriminative power of learned representations. However, efficiently leveraging hard negatives remains a challenge due to the difficulty in identifying and incorporating them without significantly increasing computational costs. To address this, we introduce SynCo (Synthetic negatives in Contrastive learning), a novel contrastive learning approach that improves model performance by generating synthetic hard negatives. Built on the MoCo framework, SynCo introduces six novel strategies for creating diverse synthetic hard negatives that can be generated on-the-fly with minimal computational overhead. SynCo achieves faster training and better representation learning, achieving a top-1 accuracy of 68.1% in ImageNet linear evaluation after only 200 epochs on pretraining, surpassing MoCo’s 67.5% with the same ResNet-50 encoder. Additionally, it transfers more effectively to detection tasks: on the PASCAL VOC, it outperforms both the supervised baseline and MoCo, achieving an $AP$ of 82.5%; on the COCO dataset, it sets a new benchmark with 40.4% $AP$ for bounding box detection and 35.4% $AP$ for instance segmentation. Our synthetic hard negative generation procedure significantly enhances the quality of visual representations learned through self-supervised contrastive learning. Code is available at https://github.com/giakoumoglou/synco.
1 Introduction
Figure 1: SynCo extends MoCo by introducing synthetic hard negatives generated on-the-fly from a memory queue. The process begins with two augmented views of an image, $\mathbf{x}_{q}$ and $\mathbf{x}_{k}$, processed by an encoder and a momentum encoder, respectively, producing feature vectors $\mathbf{q}$ and $\mathbf{k}$. The memory queue holds negative samples $\mathbf{n}_{1},\mathbf{n}_{2},\ldots$, which are concatenated with synthetic hard negatives $\mathbf{h}_{1},\mathbf{h}_{2},\ldots$ generated using the SynCo strategies. These combined negatives are used to compute the affinity matrix, which, together with the positive pair (query $\mathbf{q}$ and key $\mathbf{k}$), contributes to the InfoNCE loss calculation. This approach enhances the learning process by providing more challenging and informative negative samples.
Contrastive learning has emerged as a prominent approach in self-supervised learning, significantly advancing representation learning from unlabeled data. This technique, which discriminates between similar and dissimilar data pairs, has shown premise in visual representation tasks. Seminal works such as SimCLR [6] and MoCo [17] established instance discrimination as a pretext task. These methods generate multiple views of the same data point through augmentation, training the model to minimize the distance between positive pairs (augmented views of the same instance) while maximizing it for negative pairs (views of different instances).
Despite its efficacy, instance discrimination faces significant challenges. A key limitation is its dependence on numerous negative samples for optimal performance, often leading to increased computational costs. For example, SimCLR requires large batch sizes to provide sufficient negatives [6]. While approaches like MoCo mitigate some issues through a dynamic queue and momentum encoder [17, 8], they still confront the challenge of selecting and maintaining high-quality hard negatives—samples particularly similar to positives and thus more difficult to distinguish.
Recent studies [6, 10, 33, 39, 8, 20, 29, 3] highlight the importance of carefully crafted data augmentations in learning robust representations. These transformations likely provide more diverse, challenging copies of images, increasing the difficulty of the self-supervised task. This self-supervised task is a pretext problem (e.g., predicting image rotations [12] or solving jigsaw puzzles [27]) designed to induce learning of generalizable features without explicit labels. At the same time, data mixing techniques operating at either the pixel [46, 43] or the feature level [37] help models learn more robust features that improve both supervised and semi-supervised learning on subsequent (target) tasks.
The concept of hard negatives has been explored extensively as a means to enhance the robustness and generalization of contrastive learning models. Hard negatives are those that lie close to the decision boundary, making them crucial for fine-tuning the model’s discriminative capabilities. [20] introduced MoCHI, which integrates hard negative mixing into the MoCo framework, demonstrating significant improvements in performance. The task of efficiently creating and choosing hard negatives is important in contrastive self-supervised learning.
In this paper, we present SynCo (Synthetic negatives in Contrastive learning), a novel approach to contrastive learning that leverages synthetic hard negatives to enhance the learning process. Building on the foundations of MoCo, SynCo introduces six distinct strategies for generating synthetic hard negatives, each designed to provide diverse and challenging contrasts to the model. These strategies include: (1) Interpolated negatives; (2) Extrapolated negatives; (3) Mixup negatives; (4) Noise-injected negatives; (5) Perturbed negatives; and (6) Adversarial negatives. By incorporating these synthetic samples, SynCo aims to push the boundaries of contrastive learning, improving both the efficiency and effectiveness of the training process.
The main contributions of our work are as follows:
  1. 1.
    We introduce SynCo, a novel contrastive learning approach that improves representation learning by leveraging synthetic hard negatives. SynCo enhances the discriminative capabilities of models by generating challenging negatives on-the-fly from a memory queue, using six distinct strategies that target different aspects of the feature space to provide a comprehensive approach to hard negative generation. This process improves model performance without significant increases in computational demands, achieving faster training and better representation learning.
  2. 2.
    We empirically show improved downstream performance on ImageNet by incorporating synthetic hard negatives, demonstrating enhancements in both linear evaluation and semi-supervised learning tasks.
  3. 3.
    We show SynCo learns stronger representations by measuring their transfer learning capabilities COCO and PASCAL VOC detection, where it outperforms both the supervised baseline and MoCo.
The paper is structured as follows: Section 2 reviews related work; Section 3 explores hard negatives in contrastive learning; Section 4 introduces our synthetic hard negatives method; Section 5 presents experimental results; and Section 6 concludes the paper.
2 Related Work
2.1 Contrastive Learning
Present contrastive learning methods learn representations by instance discrimination as a pretext task, treating each image as its own class. The core idea is to pull together an anchor and a ”positive” sample in the embedding space while pushing apart the anchor from ”negative” samples [6, 21]. Positive pairs are typically generated by creating multiple views of each data point [33], such as through luminance and chrominance decomposition [32], random augmentation [6, 17], image patches [36], or student-teacher model representations [13, 5, 28]. This approach achieves state-of-the-art performance on downstream tasks. The common training objective, provided by InfoNCE [36] or its variants [6, 10, 35], maximizes mutual information [18, 2], requiring numerous negative pairs. While large batch sizes [6] address this, they’re resource-intensive. Alternatively, MoCo [17, 8], PIRL [25], and InstDis [41] use memory structures. Recent advancements explore strategies like invariance regularizers [26], dataset-derived positives [10], and unified contrastive formulas [31]. Some approaches eliminate negative samples via asymmetric Siamese structures or normalization [13, 7, 5], while others prevent model collapse through redundancy reduction [44] or regularization [4]. Despite these innovations, instance discrimination’s fine-grained nature can lead to false-negative pairs [47].
2.2 Hard Negatives
In contrastive learning, using hard negatives is crucial for improving the learning process and model performance. Hard negatives are challenging yet relevant samples that help in defining the embedding space more effectively, as they allow the model to better distinguish between similar features. The use of hard negatives involves selecting negatives that are similar to positive samples but are different enough to aid the model in learning distinctive features. Dynamically sampling these hard negatives during training prevents the model from easily minimizing the loss, which enhances its learning capabilities [17, 6]. MoCo [17] uses a dynamic queue and a momentum-based encoder update mechanism. This method keeps the negatives fresh by periodically updating the queue with new samples, which maintains the challenge for the model throughout the training period. SimCLR [6] and InfoMin [33] suggest that the difficulty of negative samples can be adjusted by varying the data augmentation techniques used during training. This makes the task progressively harder, which is beneficial for training. Building on these ideas, MoCHI [20] extends the MoCo framework by introducing hard negative mixing into the embedding space.
When integrating these approaches into embedding space generation, it is essential to consider how negative samples impact the model’s learning space. The goal is to implement a strategy that not only provides a challenge to the model but also enhances its ability to extract and differentiate features effectively. By using these methods models become more capable of handling detailed and complex tasks. This approach ensures that each negative sample significantly contributes to the training, optimizing learning outcomes and boosting overall model effectiveness.
3 Hard Negatives in Contrastive Learning
3.1 Contrastive Learning with Memory
Contrastive learning seeks to differentiate between similar and dissimilar data pairs, often treated as a dictionary look-up where representations are optimized to align positively paired data through contrastive loss in the embedding space [17]. Given an image $x$, and a distribution of image augmentation $\mathcal{T}$, we create two augmented views of the same image using the transformation $t_{q},t_{k}\sim\mathcal{T}$, i.e., $x_{q}=t_{q}(x)$ and $x_{k}=t_{k}(x)$. Two encoders, $f_{q}$ and $f_{k}$, namely the query and key encoders, generate the vectors $\mathbf{q}=f_{q}(x_{q})$ and $\mathbf{k}=f_{k}(x_{k})$ respectively. The learning objective minimizes a contrastive loss using the InfoNCE criterion [36]:
$\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q})=-\log\frac{\exp(\mathbf{q}\cdot\mathbf{k}/\tau)}{\exp(\mathbf{q}\cdot\mathbf{k}/\tau)+\sum\limits_{\mathbf{n}\in\mathcal{Q}}\exp(\mathbf{q}\cdot\mathbf{n}/\tau)}$ (1)
where $\mathbf{k}$ is $f_{k}$’s output from the same augmented image as $\mathbf{q}$, and $\mathcal{Q}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ includes outputs from different images, representing negative samples of size $K$. The temperature parameter $\tau$ adjusts scaling for the $\ell_{2}$-normalized vectors $\mathbf{q}$ and $\mathbf{k}$. The memory bank $\mathcal{Q}$ can be defined as an external memory containing every other image in the dataset [25, 32, 41], a queue of the last batches [17], or every other image in the current minibatch [6].
The log-likelihood function of the contrastive loss in Equation 1 is defined over the probability distribution created by applying a softmax function for each input/query $\mathbf{q}$. The gradient of the loss with respect to the query $\mathbf{q}$ is given by:
$\frac{\partial\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q})}{\partial\mathbf{q}}=-\frac{1}{\tau}\left((1-p_{k})\cdot\mathbf{k}-\sum_{\mathbf{n}\in\mathcal{Q}}p_{n}\cdot\mathbf{n}\right)$ (2)
where
$p_{z_{i}}=\frac{\exp(\mathbf{q}\cdot\mathbf{z_{i}}/\tau)}{\sum_{j\in Z}\exp(\mathbf{q}\cdot\mathbf{z_{j}}/\tau)}$ (3)
with $\mathbf{z_{i}}$ being a member of the set $\mathcal{Q}\cup\{\mathbf{k}\}$. We see that the contributions of the positive and negative logits to the loss are identical to the ones for a $(K+1)$-way cross-entropy classification loss, where the logit for the key corresponds to the query’s latent class [1] and all gradients are scaled by $1/\tau$.
3.2 Understanding Hard Negatives
Hard negatives are critical for contrastive learning [1, 14, 19, 24, 41, 20]. Sampling negatives from the same batch leads to a need for larger batches [6] while sampling negatives from a memory bank that contains every other image in the dataset requires the time consuming task of keeping a large memory up-to-date [25, 41]. In the latter case, a trade-off exists between the ”freshness” of the memory bank representations and the computational overhead for re-computing them as the encoder keeps changing. MoCo [17] offers a compromise between the two negative sampling extremes: it keeps a queue of the latest $K$ features from the last batches, encoded with a second key encoder that trails the (main/query) encoder with a much higher momentum. For MoCo, the key feature $\mathbf{k}$ and all features in $\mathcal{Q}$ are encoded with the key encoder.
Hardness of negatives. Figure 2 illustrates the evolving difficulty of negatives during MoCo-v2 training on ImageNet-100. We plot the top 1024 matching probabilities $p_{z_{i}}$. Initially, at epoch 0, the logits exhibit a relatively uniform distribution. However, as training progresses, an increasing number of negatives cease to contribute significantly to the loss. This suggests that the majority of memory negatives become less effective for the proxy task over time. Previous works [17, 6] have shown that larger memory or batch sizes are necessary for obtaining more challenging negatives. Our analysis confirms that many memory negatives provide diminishing benefits to the learning process as training advances.
Figure 2: Performance comparison of MoCo, MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of accuracy on the proxy task (percentage of queries where we rank the key over all negatives).
Difficulty of the proxy task. MoCo and SimCLR learn augmentation-invariant representations. We follow the hypothesis of [20] that proxy task difficulty correlates with transformation set complexity. Figure 3 shows the proxy task performance (percentage of queries ranking the key over all negatives) for MoCo and MoCo-v2 on ImageNet-100. MoCo-v2, despite lower proxy task performance due to harder augmentations, achieves better linear classification results compared to MoCo. This suggests that increasing proxy task difficulty through more complex augmentations can enhance overall performance.
Figure 3: Performance comparison of MoCo, MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of accuracy on the proxy task (percentage of queries where we rank the key over all negatives).
4 Synthetic Hard Negatives in Contrastive Learning
Based on these observations, we present an approach for generating synthetic hard negatives in the embedding space using six distinct strategies. We refer to the proposed approach as SynCo (”Synthetic negatives in Contrastive learning”).
4.1 Hard Negatives Generation
Let $\mathbf{q}$ be the query image, $\mathbf{k}$ its key, and $\mathbf{n}\in\mathcal{Q}$ represent the negative features from a memory structure of size $K$. The loss for the query is calculated using logits $\ell(z_{i})=\mathbf{q}\cdot z_{i}/\tau$ fed into a softmax function. Define $\mathcal{\hat{Q}}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ as the ordered set of all negative features, such that $\ell(\mathbf{n}_{i})>\ell(\mathbf{n}_{j})$ for all $i<j$, meaning the set of negative features is sorted by decreasing similarity to the query feature. The hardest negatives are identified by truncating the ordered set $\mathcal{\hat{Q}}$, keeping only the first $N<K$ items, denoted as $\mathcal{\hat{Q}}^{N}$.
Interpolated hard negatives (type 1). For each query $\mathbf{q}$, we propose to generate $N_{1}$ hard negative features by mixing the query $\mathbf{q}$ with a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $H^{1}=\{\mathbf{h}_{1}^{1},\mathbf{h}_{2}^{1},\ldots,\mathbf{h}_{N_{1}}^{1}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{1}\in H^{1}$ would be given by:
$\mathbf{h}_{k}^{1}=\alpha_{k}\cdot\mathbf{q}+(1-\alpha_{k})\cdot\mathbf{n}_{i},\quad\alpha_{k}\in(0,\alpha_{\text{max}})$ (4)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\alpha_{k}$ is randomly sampled from a uniform distribution in the range $(0,\alpha_{\text{max}})$. The resulting synthetic hard negatives are then normalized and added to the set of negative logits for the query. Interpolation creates a synthetic embedding that lies between the query and the negative in the embedding space. We set $\alpha_{\text{max}}=0.5$ to ensure that the query’s contribution is always smaller than the one of the negative. This is similar to the hardest negatives (type 2) of MoCHI.
Extrapolated hard negatives (type 2). For each query $\mathbf{q}$, we propose to generate $N_{2}$ hard negative features by extrapolating beyond the query embedding in the direction of the hardest negative features. Similar to the interpolated method, we use a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$. Let $H^{2}=\{\mathbf{h}_{1}^{2},\mathbf{h}_{2}^{2},\ldots,\mathbf{h}_{N_{2}}^{2}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{2}\in H^{2}$ would be given by:
$\mathbf{h}_{k}^{2}=\mathbf{n}_{i}+\beta_{k}\cdot(\mathbf{n}_{i}-\mathbf{q}),\quad\beta_{k}\in(1,\beta_{\text{max}})$ (5)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\beta_{k}$ is randomly sampled from a uniform distribution in the range $(1,\beta_{\text{max}})$. These synthetic features are also normalized and used to enhance the negative logits. Extrapolation generates a synthetic embedding that lies beyond the query embedding in the direction of the hardest negative. We choose $\beta_{\text{max}}=1.5$.
Mixup hard negatives (type 3). For each query $\mathbf{q}$, we propose to generate $N_{3}$ hard negative features by combining pairs of the $N$ hardest existing negative features in $\mathcal{\hat{Q}}^{N}$. Let $H^{3}=\{\mathbf{h}_{1}^{3},\mathbf{h}_{2}^{3},\ldots,\mathbf{h}_{N_{3}}^{3}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{3}\in H^{3}$ would be given by:
$\mathbf{h}_{k}^{3}=\gamma_{k}\cdot\mathbf{n}_{i}+(1-\gamma_{k})\cdot\mathbf{n}_{j},\quad\gamma_{k}\in(0,1)$ (6)
where $\mathbf{n}_{i},\mathbf{n}_{j}\in\mathcal{\hat{Q}}^{N}$ and $\gamma_{k}$ is randomly sampled from a uniform distribution in the range $(0,1)$. The resulting synthetic hard negatives are then normalized and added to the set of negative logits for the query. Mixup combines pairs of the hardest existing negative features to create a synthetic embedding that represents a blend of challenging cases. This is similar to the hard negatives (type 1) of MoCHI.
Noise-injected hard negatives (type 4). For each query $\mathbf{q}$, we propose to generate $N_{4}$ hard negative features by adding Gaussian noise to the hardest negative features. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{4}=\{\mathbf{h}_{1}^{4},\mathbf{h}_{2}^{4},\ldots,\mathbf{h}_{N_{4}}^{4}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{4}\in H^{4}$ would be given by:
$\mathbf{h}_{k}^{4}=\mathbf{n}_{i}+\mathcal{N}(0,\sigma^{2})$ (7)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise with standard deviation $\sigma$. The noisy negatives are normalized before being used in the loss calculation. Noise injection adds Gaussian noise to the hardest negative features, resulting in a synthetic embedding with added randomness.
Perturbed hard negatives (type 5). For each query $\mathbf{q}$, we propose to generate $N_{5}$ hard negative features by perturbing the embeddings of the hardest negative features. Given the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{5}=\{\mathbf{h}_{1}^{5},\mathbf{h}_{2}^{5},\ldots,\mathbf{h}_{N_{5}}^{5}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{5}\in H^{5}$ would be given by:
$\mathbf{h}_{k}^{5}=\mathbf{n}_{i}+\delta\cdot\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i})$ (8)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\text{sim}(\cdot,\cdot)$ is the similarity function and $\delta$ controls the perturbation magnitude. We approximate the gradients by calculating the difference in similarity scores with small perturbations applied. The perturbed embeddings are then normalized and added to the negative logits. Perturbation modifies the embeddings of the hardest negative features based on the gradient of the similarity function, creating synthetic negatives that are slightly adjusted to be more challenging for the model.
Adversarial hard negatives (type 6). For each query $\mathbf{q}$, we propose to generate $N_{6}$ hard negative features by applying adversarial perturbations to the hardest negative features to maximize their similarity to the query embeddings. Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{6}=\{\mathbf{h}_{1}^{6},\mathbf{h}_{2}^{6},\ldots,\mathbf{h}_{N_{6}}^{6}\}$ be the set of synthetic negatives to be generated. Then a synthetic point $\mathbf{h}_{k}^{6}\in H^{6}$ would be given by:
$\mathbf{h}_{k}^{6}=\mathbf{n}_{i}+\eta\cdot\text{sign}(\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i}))$ (9)
where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\eta$ controls the perturbation magnitude. The perturbed embeddings are normalized and added to the negative logits. Adversarial hard negatives apply adversarial perturbations to the hardest negative features, specifically altering them to maximize their similarity to the query embeddings, thereby producing the most challenging contrasts.
4.2 Discussion on Synthetic Hard Negatives
Intuition of synthetic hard negatives. Each SynCo strategy enhances model generalization through challenging contrasts. Type 1 interpolates between query and hard negatives, increasing sample diversity throughout training. Type 2 extrapolates beyond the query, pushing embedding space boundaries and improving robustness to difficult contrasts. Type 3 combines pairs of hard negatives, encouraging more generalized and robust feature learning. Type 4 injects Gaussian noise, promoting invariance to minor feature fluctuations and enhancing generalization. Type 5 modifies embeddings based on similarity gradients, refining discriminatory power by directing the model towards harder negatives. Type 6 applies adversarial perturbations, creating the most challenging contrasts to distinguish deceptively similar samples. A toy example of the proposed synthetic hard negative generation is presented in Figure 4. The latter figure visualizes the different types of synthetic hard negatives generated from a query point and three original negative points in a 2D embedding space. Dotted lines connect the query to extrapolated negatives and between pairs of mixup negatives, illustrating the relationships used to generate these synthetic points.
Figure 4: A toy example illustrating the six types of synthetic hard negatives generated by SynCo from a query point and three original negative points in a 2D embedding space.
Is the proxy task more difficult? Figure 3 illustrates the proxy task performance for various configurations of SynCo. We observe that incorporating synthetic negatives leads to faster learning and improved performance. Each type of synthetic negative accelerates learning compared to the MoCo-v2 baseline, with the full SynCo configuration showing the most significant improvement (see Figure 6(a)) and the lowest final proxy task performance. This indicates that SynCo presents the most challenging proxy task. The lower performance when including synthetic negatives suggests that $\max\ell(\mathbf{h}_{k}^{i})>\max\ell(\mathbf{n}_{j})$, $\mathbf{h}_{k}^{i}\in H^{i}$, $\mathbf{n}_{j}\in\tilde{\mathcal{Q}}_{N}$. Through SynCo, we modulate proxy task difficulty via synthetic negatives, pushing the model to learn more robust features.
Evaluating the usage of the embedding space. Wang and Isola [38] presented two losses/metrics for assessing contrastive learning representations. The first metric assesses alignment by measuring the absolute distance between representations that share the same label on a hypersphere. The second metric evaluates uniformity by calculating the logarithm of the average pairwise Gaussian potential among all embeddings. In Figure 5, we present these metrics for various models, based on re-implemented methods and experiments conducted using features from all images in the ImageNet-100 validation set. Our findings demonstrate that SynCo significantly enhances the uniformity of representations compared to MoCo-v2 and MoCHI, indicating its superior ability to utilize the embedding space in the proxy task. Furthermore, the incorporation of synthetic negatives (types 1 to 6) leads to improved alignment. These results suggest that SynCo’s approach to synthetic negative generation and contrastive learning yields more effective and well-distributed feature representations.
Figure 5: Performance comparison of MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of alignment and uniformity metrics. The x/y axes correspond to $-\mathcal{L}_{\text{uniform}}$ and $-\mathcal{L}_{\text{align}}$, respectively. The best-performing model lies towards the top-right of the plot. We use $K=65$k.
5 Experiments
5.1 Implementation Details
We learn representations on the common ImageNet ILSVRC-2012 [9] and its smaller ImageNet-100 subset [32] using a ResNet-50 [15] encoder. All runs of SynCo are based on MoCo-v2 [8]. We developed our approach on top of the official public implementation of MoCo and reproduced it on our setup. We run all experiments on 4 NVIDIA A40 GPUs. For linear classification on ImageNet, we follow the common protocol and report results on the validation set. We report performance after learning linear classifiers for 100 epochs, with an initial learning rate of $30.0$ ($10.0$ for ImageNet-100), a batch size of 256 and a cosine learning rate schedule. For training, unless stated otherwise, we use $K=65$k ($K=16$k for ImageNet-100). For SynCo, we also have a warm-up of 10 epochs, i.e. for the first epochs we do not synthesize hard negatives. We set SynCo’s hyperparameters $\sigma$, $\delta$, and $\eta$ to 0.01. For hard negative generation, we use the top $N=1024$ hardest negatives, with $N_{1}=N_{2}=N_{3}=256$ and $N_{4}=N_{5}=N_{6}=64$. For ImageNet, we report accuracy for a single-crop testing. For object detection on PASCAL VOC [11] we follow [17] and fine-tune a Faster R-CNN [30], R50-C4 on trainval07+12 and test on test2007. For COCO object detection, we follow [17] and fine-tune a Mask R-CNN [16], R50-C4 on train2017 and test on val2017. We use the open-source Detectron2 [40] code and report the common $AP$, $AP_{50}$ and $AP_{75}$ metrics. Similar to [17], we do not perform hyperparameter tuning for the object detection task. Detailed implementation details are provided in the supplementary material.
5.2 Linear Evaluation on ImageNet
We evaluate the SynCo representation by training a linear classifier on top of the frozen features that were pretrained on ImageNet, following the procedure described in [23, 22, 6, 13, 36]. Full details of the linear evaluation protocol are described in the supplementary material. With a standard ResNet-50 and for 200 epoch pretraining, SynCo obtains $67.9\%\pm 0.16\%$ top-1 accuracy and $88.0\%\pm 0.05\%$ top-5 accuracy (Table 1). Specifically, SynCo achieves a +0.4% top-1 accuracy improvement over MoCo-v2 and a +1.0% improvement over MoCHI with 200 epochs of pretraining, leveraging synthetic hard negatives. While MoCHI, which employs hard negatives, achieves lower performance than MoCo-v2, our method, which generates synthetic hard negatives, not only avoids this drop in performance but actually improves it. SynCo also surpasses the state-of-the-art methods SimCLR and SimSiam. When training for 800 epochs, SynCo obtains 70.6% top-1 accuracy (89.8% top-5 accuracy) (Table 2). This represents a +1.9% improvement in top-1 accuracy over MoCHI and a -0.5% decrease compared to MoCo-v2 under the same 800 epochs pretraining settings. These results show that SynCo achieves higher accuracy than its competitors with only a minor computational overhead for generating synthetic hard negatives, while also enabling faster training convergence.
Table 1: Top-1 and top-5 accuracies (in %) under linear evaluation on ImageNet ILSVRC-2012 with 200 epochs of pretraining using ResNet-50. Results for SynCo are averaged over 3 runs.
Method Top-1 Top-5
Supervised 76.5 -
PIRL 63.6 -
LA 60.2 -
InfoMin 70.1 89.4
SimSiam 68.1 -
SimCLR 69.3 -
MoCo 60.7 -
MoCo-v2 67.5 90.1
MoCHI 66.9 -
SynCo (ours) 67.9 88.0
Table 2: Top-1 and top-5 accuracies (in %) under linear evaluation on ImageNet ILSVRC-2012 for models trained with extended epochs using ResNet-50. Results for SynCo are based on 1 run.
Method Epochs Top-1 Top-5
PIRL 800 63.6 -
InfoMin 800 73.0 91.1
BYOL 1000 74.3 91.6
SimSiam 800 68.1 -
SimCLR 1000 69.3 -
Barlow Twins 1000 73.2 91.0
MoCo-v2 800 71.1 90.1
MoCHI 800 68.7 -
SynCo (ours) 800 70.6 89.8
5.3 Semi-supervised Training on ImageNet
We evaluate the performance obtained when fine-tuning SynCo representation, pretrained for 800 epochs, on a classification task with a small subset of ImageNet’s train set, this time using label information. We follow the semi-supervised protocol of [23, 45, 6, 13, 36], and use the same fixed splits of respectively 1% and 10% of ImageNet labeled training data as in [6]. Our complete semi-supervised learning setup is detailed in the supplementary material. Despite not being specifically designed for this task, SynCo shows promising results (Table 3). With 1% labeled data, SynCo achieves 50.8% $\pm$ 0.21% top-1 accuracy and 77.5% $\pm$ 0.12% top-5 accuracy. With 10% labeled data, SynCo reaches 66.6% $\pm$ 0.19% top-1 accuracy and 88.0% $\pm$ 0.10% top-5 accuracy. We observe that despite a lower volume of labeled data, SynCo competes well with established semi-supervised methods like SimCLR, BYOL, and Barlow Twins.
Table 3: Semi-supervised learning on ImageNet ILSVRC-2012 with 1% and 10% training examples using ResNet-50. Results for SynCo are averaged over 3 runs.
Method Epochs Top-1 Top-5
1% 10% 1% 10%
Supervised 25.4 56.4 48.4 80.4
InstDis 200 - - 39.2 77.4
PIRL 800 30.7 60.4 57.2 83.8
SimCLR 1000 48.3 65.6 75.5 87.8
Barlow Twins 1000 55.0 69.7 79.2 89.3
BYOL 1000 53.2 68.8 78.4 89.0
SwAV 800 53.9 70.2 78.5 89.9
SynCo (ours) 800 50.8 66.6 77.5 88.0
5.4 Transferring to Detection
We evaluate the SynCo representation, pretrained for 200 epochs, by applying it to the detection task. We follow the protocol of [17], and as used in [20] for a fair comparison. Detailed configurations for object detection experiments are provided in in the supplementary material. Table 4 presents the results of object detection on PASCAL VOC and COCO datasets. SynCo displays consistent gains over both the supervised baseline and MoCo-v2. Specifically, SynCo, surpasses MoCo-v2 and performs on par with MoCHI in the PASCAL VOC detection task. On the more challenging COCO dataset, SynCo achieves new state-of-the-art results in both bounding box detection and instance segmentation. Additional results comparing SynCo and MoCo-v2 at different training epochs are presented in the supplementary material.
Table 4: Results for object detection on PASCAL VOC and COCO. All models use a ResNet-50 backbone. We report AP, AP50, and AP75, which are standard COCO metrics. $bb$ denotes bounding box detection, and $msk$ denotes instance segmentation. The values in bold indicate the maximum of each column. Results for SynCo are averaged over 3 runs.
Method VOC07+12 detection COCO detection COCO segmentation
$AP$ $AP_{50}$ $AP_{75}$ $AP^{bb}$ $AP^{bb}_{50}$ $AP^{bb}_{75}$ $AP^{msk}$ $AP^{msk}_{50}$ $AP^{msk}_{75}$
Random init 16.8 35.9 13.0 31.0 49.5 33.2 28.5 46.8 30.4
Supervised 53.5 81.3 58.8 38.2 58.2 41.6 33.3 54.7 35.2
InstDis - - - 37.4 57.6 40.6 34.1 54.6 36.4
PIRL 55.5 81.0 61.3 38.5 57.6 41.2 34.0 54.6 36.2
InfoMin 57.6 82.7 64.6 39.0 58.5 42.0 34.1 55.2 36.3
SwAV 56.1 82.6 62.7 38.2 58.2 41.3 33.8 55.2 35.9
SimSiam 57.0 82.4 63.7 39.2 59.3 42.1 34.4 56.0 36.7
Barlow Twins 56.8 82.6 63.4 39.2 59.0 42.5 34.3 56.0 36.5
MoCo 55.5 81.5 61.3 38.5 58.3 41.6 33.6 54.8 35.6
MoCo-v2 57.0 82.4 63.6 39.0 58.6 41.9 34.2 55.4 36.2
MoCHI 57.5 82.7 64.4 39.2 58.9 42.4 34.4 55.5 36.6
SynCo (ours) 57.2 82.5 63.9 40.4 60.1 43.8 35.4 57.0 37.4
5.5 Ablation Study
Figure 6: Ablation studies on ImageNet-100 in terms of Top-1 and Top-5 accuracies (in %). (a) Performance comparison of different types of hard negatives. (b) Performance comparison of different values for $\sigma$, $\delta$, and $\eta$ on SynCo’s type 4, type 5, and type 6 hard negatives, respectively. (c) Comparison of SynCo and MoCo-v2 across different queue sizes. For all ablations we use $K=65$k.
We perform ablations studies on ImageNet-100. The results of our ablations are presented in Figure 6. Our findings consistently demonstrate that various SynCo configurations outperform the MoCo-v2 baseline. Additional ablation studies and analyses are presented in the supplementary material.
Type of hard negative. We evaluate the impact of each synthetic hard negative type on pretraining. For this, we select the top $N=1024$ hardest negatives and generate $N_{i}=256$, $i=1,2,\ldots,6$ negatives. We train SynCo without hard negatives (equivalent to MoCo-v2) for 100 epochs and measure top-1 and top-5 accuracy. Subsequently, we train SynCo using each type of hard negative individually, and then using all six types in combination. The results of these ablations are presented in Figure 6(a). We see that every SynCo configuration outperform the MoCo-v2 baseline.
Parameter sensitivity analysis. We conducted ablations on the parameters $\sigma$, $\delta$, and $\eta$ of SynCo’s type 4, type 5, and type 6 negatives, respectively. The results, presented in Figure 6(b), show that varying these parameters does not lead to significant differences in performance. This suggests that SynCo is robust across a wide range of values for $\sigma$, $\delta$, $\eta$.
Queue size. We investigate the effect of queue size $\mathcal{Q}$ on performance. We train SynCo and MoCo-v2 with reduced queue sizes. Our results, presented in Figure 6(c), reveal that SynCo performs comparably to MoCo-v2 across various queue sizes. With smaller queues, SynCo underperforms compared to MoCo-v2. This can be attributed to the fact that the total generated synthetic negatives are too hard for the task and harm performance, a finding that is also observed in [20]. However, as the queue increases, SynCo performs on par with MoCo-v2. At the largest queue size tested, SynCo outperforms MoCo-v2.
6 Conclusion
In this paper, we built upon the observation that harder negatives are needed for more effective self-supervised contrastive learning. Based on that, we presented SynCo, a synthetic hard negative generation approach that improves the quality of representations learned in an unsupervised way, offering better transfer learning performance and utilization of the embedding space. We showed that our method learns generalizable representations, which is important considering the high compute cost of self-supervised learning. We found that multiple SynCo configurations provide considerable gains, and that using synthetic hard negatives consistently improves transfer learning performance.
While our experiments primarily utilized the MoCo framework, the proposed hard negative generation strategies are general and can be applied to any contrastive learning method that benefits from hard negatives, including SimCLR[6], CPC [36], PIRL [25], and others [42, 39, 21, 10, 34]. These methods, which employ the InfoNCE loss function (or its variants [6, 10]) and instance discrimination as the pretext task, can benefit from the enhanced hard negative generation strategies proposed by SynCo. The introduction of synthetic hard negatives can provide these methods with more challenging and informative contrasts, potentially leading to better feature representations. Also, the applicability of SynCo extends beyond visual representation learning. The concept of hard negatives can be applied to various domains such as natural language processing, audio processing, and other areas where contrastive learning is relevant.

### v3

@@
 Abstract
-Contrastive learning has become a dominant approach in self-supervised visual representation learning, with hard negatives—samples that closely resemble the anchor—being key to enhancing the discriminative power of learned representations.
-However, efficiently leveraging hard negatives remains a challenge due to the difficulty in identifying and incorporating them without significantly increasing computational costs.
-To address this, we introduce SynCo (Synthetic negatives in Contrastive learning), a novel contrastive learning approach that improves model performance by generating synthetic hard negatives.
-Built on the MoCo framework, SynCo introduces six novel strategies for creating diverse synthetic hard negatives that can be generated on-the-fly with minimal computational overhead.
-SynCo achieves faster training and better representation learning, achieving a top-1 accuracy of 68.1% in ImageNet linear evaluation after only 200 epochs on pretraining, surpassing MoCo’s 67.5% with the same ResNet-50 encoder.
-Additionally, it transfers more effectively to detection tasks: on the PASCAL VOC, it outperforms both the supervised baseline and MoCo, achieving an $AP$ of 82.5%; on the COCO dataset, it sets a new benchmark with 40.4% $AP$ for bounding box detection and 35.4% $AP$ for instance segmentation.
-Our synthetic hard negative generation procedure significantly enhances the quality of visual representations learned through self-supervised contrastive learning.
+Contrastive learning has become a dominant approach in self-supervised visual representation learning.
+Hard negatives - samples closely resembling the anchor - are key to enhancing learned representations’ discriminative power.
+However, efficiently leveraging hard negatives remains challenging.
+We introduce SynCo (Synthetic negatives in Contrastive learning), a novel approach that improves model performance by generating synthetic hard negatives on the representation space.
+Building on the MoCo framework, SynCo introduces six strategies for creating diverse synthetic hard negatives on-the-fly with minimal computational overhead.
+SynCo achieves faster training and better representation learning, reaching 67.9% top-1 accuracy on ImageNet ILSVRC-201 linear evaluation after 200 pretraining epochs, surpassing MoCo’s 67.5% using the same ResNet-50 encoder.
+It also transfers more effectively to detection tasks: on PASCAL VOC, it outperforms both the supervised baseline and MoCo with 82.6% AP; on COCO, it sets new benchmarks with 41.0% AP for bounding box detection and 35.7% AP for instance segmentation.
+Our synthetic hard negative generation approach significantly enhances visual representations learned through self-supervised contrastive learning.
 Code is available at https://github.com/giakoumoglou/synco.
@@
 These combined negatives are used to compute the affinity matrix, which, together with the positive pair (query $\mathbf{q}$ and key $\mathbf{k}$), contributes to the InfoNCE loss calculation.
-This approach enhances the learning process by providing more challenging and informative negative samples.
 Contrastive learning has emerged as a prominent approach in self-supervised learning, significantly advancing representation learning from unlabeled data.
@@
 These methods generate multiple views of the same data point through augmentation, training the model to minimize the distance between positive pairs (augmented views of the same instance) while maximizing it for negative pairs (views of different instances).
-Despite its efficacy, instance discrimination faces significant challenges.
-A key limitation is its dependence on numerous negative samples for optimal performance, often leading to increased computational costs.
-For example, SimCLR requires large batch sizes to provide sufficient negatives [6].
-While approaches like MoCo mitigate some issues through a dynamic queue and momentum encoder [17, 8], they still confront the challenge of selecting and maintaining high-quality hard negatives—samples particularly similar to positives and thus more difficult to distinguish.
-Recent studies [6, 10, 33, 39, 8, 20, 29, 3] highlight the importance of carefully crafted data augmentations in learning robust representations.
+Despite its effectiveness, instance discrimination faces challenges.
+A key limitation is the need for numerous negative samples, often leading to increased computational costs.
+For example, SimCLR requires large batch sizes for sufficient negatives [6].
+While approaches like MoCo address some issues through dynamic queues and momentum encoders [20, 8], they still face challenges in selecting and maintaining high-quality hard negatives.
+Recent studies have highlighted the importance of carefully crafted data augmentations in learning robust representations [6, 11, 39, 45, 8, 23, 34, 3].
 These transformations likely provide more diverse, challenging copies of images, increasing the difficulty of the self-supervised task.
 This self-supervised task is a pretext problem (e.g., predicting image rotations [13] or solving jigsaw puzzles [32]) designed to induce learning of generalizable features without explicit labels.
-At the same time, data mixing techniques operating at either the pixel [46, 43] or the feature level [37] help models learn more robust features that improve both supervised and semi-supervised learning on subsequent (target) tasks.
-The concept of hard negatives has been explored extensively as a means to enhance the robustness and generalization of contrastive learning models.
-Hard negatives are those that lie close to the decision boundary, making them crucial for fine-tuning the model’s discriminative capabilities.
-[20] introduced MoCHI, which integrates hard negative mixing into the MoCo framework, demonstrating significant improvements in performance.
-The task of efficiently creating and choosing hard negatives is important in contrastive self-supervised learning.
+Moreover, techniques that combine data at the pixel level [53, 50] or at the feature level [43] have proven effective in helping models learn more resilient features, leading to improvements in both fully supervised and semi-supervised tasks.
+The concept of challenging negative samples has been explored as a way to enhance contrastive learning models.
+These samples, which lie close to the decision boundary, are crucial for refining the model’s discriminative abilities.
+Recent work like MoCHI [23] has shown improvements by incorporating harder negatives.
+Our work builds on this idea by proposing novel strategies for synthetic hard negative generation.
 In this paper, we present SynCo (Synthetic negatives in Contrastive learning), a novel approach to contrastive learning that leverages synthetic hard negatives to enhance the learning process.
 Building on the foundations of MoCo, SynCo introduces six distinct strategies for generating synthetic hard negatives, each designed to provide diverse and challenging contrasts to the model.
-These strategies include: (1) Interpolated negatives; (2) Extrapolated negatives; (3) Mixup negatives; (4) Noise-injected negatives; (5) Perturbed negatives; and (6) Adversarial negatives.
+These strategies include: interpolated negatives; extrapolated negatives; mixup negatives; noise-injected negatives; perturbed negatives; and adversarial negatives.
 By incorporating these synthetic samples, SynCo aims to push the boundaries of contrastive learning, improving both the efficiency and effectiveness of the training process.
@@
 2.1 Contrastive Learning
-Present contrastive learning methods learn representations by instance discrimination as a pretext task, treating each image as its own class.
-The core idea is to pull together an anchor and a ”positive” sample in the embedding space while pushing apart the anchor from ”negative” samples [6, 21].
-Positive pairs are typically generated by creating multiple views of each data point [33], such as through luminance and chrominance decomposition [32], random augmentation [6, 17], image patches [36], or student-teacher model representations [13, 5, 28].
-This approach achieves state-of-the-art performance on downstream tasks.
-The common training objective, provided by InfoNCE [36] or its variants [6, 10, 35], maximizes mutual information [18, 2], requiring numerous negative pairs.
-While large batch sizes [6] address this, they’re resource-intensive.
-Alternatively, MoCo [17, 8], PIRL [25], and InstDis [41] use memory structures.
-Recent advancements explore strategies like invariance regularizers [26], dataset-derived positives [10], and unified contrastive formulas [31].
-Some approaches eliminate negative samples via asymmetric Siamese structures or normalization [13, 7, 5], while others prevent model collapse through redundancy reduction [44] or regularization [4].
-Despite these innovations, instance discrimination’s fine-grained nature can lead to false-negative pairs [47].
+Recent contrastive learning methods focus on instance discrimination as a pretext task, treating each image as its own class.
+The core principle involves bringing an anchor and a ”positive” sample closer in the representation space while pushing the anchor away from ”negative” samples [6, 24].
+Positive pairs are typically created through multiple views of each data point [39], using techniques such as color decomposition [38], random augmentation [6, 20], image patches [42], or student-teacher model representations [14, 5, 33].
+The common training objective, based on InfoNCE [42] or its variants [6, 11, 41], aims to maximize mutual information [21, 2], necessitating numerous negative pairs.
+While some approaches use large batch sizes [6] to address this, others like MoCo [20, 8], PIRL [30], and InstDis [47] employ memory structures.
+Recent advancements explore strategies such as invariance regularizers [31], dataset-derived positives [11], and unified contrastive formulas [37].
+Some methods eliminate negative samples through asymmetric Siamese structures or normalization [14, 7, 5], while others prevent model collapse via redundancy reduction [51] or regularization [4].
+Despite these innovations, the fine-grained nature of instance discrimination can lead to false-negative pairs [55].
 2.2 Hard Negatives
-In contrastive learning, using hard negatives is crucial for improving the learning process and model performance.
-Hard negatives are challenging yet relevant samples that help in defining the embedding space more effectively, as they allow the model to better distinguish between similar features.
-The use of hard negatives involves selecting negatives that are similar to positive samples but are different enough to aid the model in learning distinctive features.
-Dynamically sampling these hard negatives during training prevents the model from easily minimizing the loss, which enhances its learning capabilities [17, 6].
-MoCo [17] uses a dynamic queue and a momentum-based encoder update mechanism.
-This method keeps the negatives fresh by periodically updating the queue with new samples, which maintains the challenge for the model throughout the training period.
-SimCLR [6] and InfoMin [33] suggest that the difficulty of negative samples can be adjusted by varying the data augmentation techniques used during training.
-This makes the task progressively harder, which is beneficial for training.
-Building on these ideas, MoCHI [20] extends the MoCo framework by introducing hard negative mixing into the embedding space.
-When integrating these approaches into embedding space generation, it is essential to consider how negative samples impact the model’s learning space.
-The goal is to implement a strategy that not only provides a challenge to the model but also enhances its ability to extract and differentiate features effectively.
-By using these methods models become more capable of handling detailed and complex tasks.
-This approach ensures that each negative sample significantly contributes to the training, optimizing learning outcomes and boosting overall model effectiveness.
-3 Hard Negatives in Contrastive Learning
-3.1 Contrastive Learning with Memory
-Contrastive learning seeks to differentiate between similar and dissimilar data pairs, often treated as a dictionary look-up where representations are optimized to align positively paired data through contrastive loss in the embedding space [17].
+Hard negatives are critical in contrastive learning as they improve the quality of visual representations by helping to define the representation space more effectively.
+These challenging yet relevant samples are harder to distinguish from the anchor point, enabling the model to better differentiate between similar features.
+The use of hard negatives involves selecting samples that are similar to positive samples but different enough to aid in learning distinctive features.
+Dynamic sampling of hard negatives during training prevents the model from easily minimizing the loss, enhancing its learning capabilities [20, 6].
+Various approaches have been proposed to leverage hard negatives effectively.
+For instance, MoCo [20] utilizes a dynamic queue and momentum-based encoder updates to maintain fresh and challenging negatives throughout training.
+Other methods, such as SimCLR [6] and InfoMin [39], suggest adjusting the difficulty of negative samples by varying data augmentation techniques.
+This progressive increase in task difficulty benefits the training process.
+Building on these ideas, MoCHI [23] has explored integrating hard negative mixing into existing frameworks to further improve performance.
+When incorporating these approaches into representation space generation, it’s essential to consider how negative samples impact the model’s learning space.
+The goal is to implement a strategy that not only challenges the model but also enhances its ability to extract and differentiate features effectively.
+By employing these methods, models become more adept at handling detailed and complex tasks, ensuring each negative sample significantly contributes to optimizing learning outcomes and boosting overall model effectiveness.
+2.3 Synthetic Features
+Synthetic feature generation is a widely used method to enhance deep learning models, especially in cases with limited labeled data.
+By adding synthetic features to the representation space, models improve in generalization and performance.
+Methods have generated features for unseen classes using generative models [17, 48, 36], while others have integrated these into self-supervised and contrastive learning frameworks [27, 54].
+Models that combine generation with representation spaces have also shown success in zero-shot learning [16].
+In contrast, our approach directly generates synthetic hard negatives in contrastive learning, improving representation without additional generative models.
+3 Preliminaries
+3.1 Contrastive Learning
+Contrastive learning seeks to differentiate between similar and dissimilar data pairs, often treated as a dictionary look-up where representations are optimized to align positively paired data through contrastive loss in the representation space [20].
 Given an image $x$, and a distribution of image augmentation $\mathcal{T}$, we create two augmented views of the same image using the transformation $t_{q},t_{k}\sim\mathcal{T}$, i.e., $x_{q}=t_{q}(x)$ and $x_{k}=t_{k}(x)$.
@@
 The learning objective minimizes a contrastive loss using the InfoNCE criterion [42]:
-$\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q})=-\log\frac{\exp(\mathbf{q}\cdot\mathbf{k}/\tau)}{\exp(\mathbf{q}\cdot\mathbf{k}/\tau)+\sum\limits_{\mathbf{n}\in\mathcal{Q}}\exp(\mathbf{q}\cdot\mathbf{n}/\tau)}$ (1)
-where $\mathbf{k}$ is $f_{k}$’s output from the same augmented image as $\mathbf{q}$, and $\mathcal{Q}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ includes outputs from different images, representing negative samples of size $K$.
+$\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q})=-\log\frac{\exp(\mathbf{q}^{T}\cdot\mathbf{k}/\tau)}{\exp(\mathbf{q}^{T}\cdot\mathbf{k}/\tau)+\sum\limits_{\mathbf{n}\in\mathcal{Q}}\exp(\mathbf{q}^{T}\cdot\mathbf{n}/\tau)}$ (1)
+Here, $\mathbf{k}$ is $f_{k}$’s output from the same augmented image as $\mathbf{q}$, and $\mathcal{Q}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ includes outputs from different images, representing negative samples of size $K$.
 The temperature parameter $\tau$ adjusts scaling for the $\ell_{2}$-normalized vectors $\mathbf{q}$ and $\mathbf{k}$.
-The memory bank $\mathcal{Q}$ can be defined as an external memory containing every other image in the dataset [25, 32, 41], a queue of the last batches [17], or every other image in the current minibatch [6].
-The log-likelihood function of the contrastive loss in Equation 1 is defined over the probability distribution created by applying a softmax function for each input/query $\mathbf{q}$.
-The gradient of the loss with respect to the query $\mathbf{q}$ is given by:
+The key encoder $f_{k}$ can be updated in two ways.
+In the synchronized update approach, $f_{k}$ is updated synchronously with $f_{q}$, maintaining identical weights throughout training [6].
+Alternatively, a momentum update scheme can be employed, where $f_{k}$ is updated using the equation: $\theta_{k}\leftarrow m\cdot\theta_{k}+(1-m)\cdot\theta_{q}$ [20].
+Here, $\theta_{k}$ and $\theta_{q}$ are the parameters of $f_{k}$ and $f_{q}$ respectively, and $m\in[0,1]$ is the momentum coefficient.
+This momentum approach allows $f_{k}$ to evolve more slowly, providing more consistent negative samples over time and potentially stabilizing the learning process.
+The memory bank $\mathcal{Q}$ can be defined in various ways, such as an external memory of all dataset images [30, 38, 47], a queue of recent batches [20], or the current minibatch [6].
+The gradient of the contrastive loss in Equation 1 with respect to the query $\mathbf{q}$ is given by:
 $\frac{\partial\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q})}{\partial\mathbf{q}}=-\frac{1}{\tau}\left((1-p_{k})\cdot\mathbf{k}-\sum_{\mathbf{n}\in\mathcal{Q}}p_{n}\cdot\mathbf{n}\right)$ (2)
 where
-$p_{z_{i}}=\frac{\exp(\mathbf{q}\cdot\mathbf{z_{i}}/\tau)}{\sum_{j\in Z}\exp(\mathbf{q}\cdot\mathbf{z_{j}}/\tau)}$ (3)
+$p_{z_{i}}=\frac{\exp(\mathbf{q}^{T}\cdot\mathbf{z_{i}}/\tau)}{\sum_{j\in Z}\exp(\mathbf{q}^{T}\cdot\mathbf{z_{j}}/\tau)}$ (3)
 with $\mathbf{z_{i}}$ being a member of the set $\mathcal{Q}\cup\{\mathbf{k}\}$.
-We see that the contributions of the positive and negative logits to the loss are identical to the ones for a $(K+1)$-way cross-entropy classification loss, where the logit for the key corresponds to the query’s latent class [1] and all gradients are scaled by $1/\tau$.
+The positive and negative logits contribute to the loss similarly to a $(K+1)$-way cross-entropy classification, with the key logit representing the query’s latent class [1].
 3.2 Understanding Hard Negatives
-Hard negatives are critical for contrastive learning [1, 14, 19, 24, 41, 20].
-Sampling negatives from the same batch leads to a need for larger batches [6] while sampling negatives from a memory bank that contains every other image in the dataset requires the time consuming task of keeping a large memory up-to-date [25, 41].
-In the latter case, a trade-off exists between the ”freshness” of the memory bank representations and the computational overhead for re-computing them as the encoder keeps changing.
-MoCo [17] offers a compromise between the two negative sampling extremes: it keeps a queue of the latest $K$ features from the last batches, encoded with a second key encoder that trails the (main/query) encoder with a much higher momentum.
-For MoCo, the key feature $\mathbf{k}$ and all features in $\mathcal{Q}$ are encoded with the key encoder.
+The effectiveness of contrastive learning approaches hinges critically on the utilization of hard negatives [1, 15, 22, 29, 47, 23].
+Current approaches face significant challenges in efficiently leveraging these hard negatives.
+Sampling from within the same batch necessitates larger batch sizes [6, 9], potentially straining computational resources.
+Conversely, maintaining a memory bank containing representations of the entire dataset incurs substantial computational overhead in keeping the memory up-to-date [30, 47, 20, 8].
+These limitations underscore the need for more efficient strategies to generate and utilize hard negatives in contrastive learning frameworks.
 Hardness of negatives.
-Figure 2 illustrates the evolving difficulty of negatives during MoCo-v2 training on ImageNet-100.
-We plot the top 1024 matching probabilities $p_{z_{i}}$.
-Initially, at epoch 0, the logits exhibit a relatively uniform distribution.
-However, as training progresses, an increasing number of negatives cease to contribute significantly to the loss.
-This suggests that the majority of memory negatives become less effective for the proxy task over time.
-Previous works [17, 6] have shown that larger memory or batch sizes are necessary for obtaining more challenging negatives.
-Our analysis confirms that many memory negatives provide diminishing benefits to the learning process as training advances.
-Figure 2: Performance comparison of MoCo, MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of accuracy on the proxy task (percentage of queries where we rank the key over all negatives).
+The ”hardness” of negative samples, defined by their similarity to positive samples in the representation space, determines how challenging they are for the model to differentiate, directly impacting the effectiveness of the contrastive learning process.
+Figure 2 illustrates the evolution of negative sample hardness during MoCo-v2 training on ImageNet-100.
+The plot depicts the top 1024 matching probabilities $p_{z_{i}}$ across different training epochs.
+Initially, the distribution of these probabilities is relatively uniform.
+However, as training progresses, a clear trend emerges: fewer negatives contribute significantly to the loss function.
+This observation suggests that the model rapidly learns to distinguish most negatives, leaving only a small subset that remains challenging.
+Such a phenomenon underscores the importance of maintaining a diverse pool of hard negatives throughout the training process to sustain effective learning [23].
+Figure 2: Histogram of the top 1024 matching probabilities $p_{z_{i}}$, $z_{i}\in\mathcal{Q}$ for MoCo-v2, over various epochs.
+Logits are organized in descending order, and each line indicates the mean matching probability across all queries [23].
 Difficulty of the proxy task.
-MoCo and SimCLR learn augmentation-invariant representations.
-We follow the hypothesis of [20] that proxy task difficulty correlates with transformation set complexity.
-Figure 3 shows the proxy task performance (percentage of queries ranking the key over all negatives) for MoCo and MoCo-v2 on ImageNet-100.
-MoCo-v2, despite lower proxy task performance due to harder augmentations, achieves better linear classification results compared to MoCo.
-This suggests that increasing proxy task difficulty through more complex augmentations can enhance overall performance.
-Figure 3: Performance comparison of MoCo, MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of accuracy on the proxy task (percentage of queries where we rank the key over all negatives).
+The difficulty of the proxy task in contrastive learning, typically defined by the self-supervised objective, significantly influences the quality of learned representations.
+Figure 3 compares the proxy task performance of MoCo and MoCo-v2 on ImageNet-100, measured by the percentage of queries where the key ranks above all negatives.
+Notably, MoCo-v2, which employs more aggressive augmentations, exhibits lower proxy task performance compared to MoCo, indicating a more challenging learning objective.
+Paradoxically, this increased difficulty correlates with improved performance on downstream tasks such as linear classification [23].
+This counterintuitive relationship between proxy task difficulty and downstream performance suggests that more challenging self-supervised objectives can lead to the learning of more robust and transferable representations, motivating the development of strategies to dynamically modulate task difficulty during training.
+Figure 3: Performance comparison of MoCo, MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of accuracy on the proxy task (percentage of queries where the key is ranked higher than all negatives).
 4 Synthetic Hard Negatives in Contrastive Learning
-Based on these observations, we present an approach for generating synthetic hard negatives in the embedding space using six distinct strategies.
+Based on these observations, we present an approach for generating synthetic hard negatives in the representation space using six distinct strategies.
 We refer to the proposed approach as SynCo (”Synthetic negatives in Contrastive learning”).
 4.1 Hard Negatives Generation
-Let $\mathbf{q}$ be the query image, $\mathbf{k}$ its key, and $\mathbf{n}\in\mathcal{Q}$ represent the negative features from a memory structure of size $K$.
-The loss for the query is calculated using logits $\ell(z_{i})=\mathbf{q}\cdot z_{i}/\tau$ fed into a softmax function.
-Define $\mathcal{\hat{Q}}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ as the ordered set of all negative features, such that $\ell(\mathbf{n}_{i})>\ell(\mathbf{n}_{j})$ for all $i<j$, meaning the set of negative features is sorted by decreasing similarity to the query feature.
-The hardest negatives are identified by truncating the ordered set $\mathcal{\hat{Q}}$, keeping only the first $N<K$ items, denoted as $\mathcal{\hat{Q}}^{N}$.
+Let $\mathbf{q}$ represent the query image, $\mathbf{k}$ its corresponding key, and $\mathbf{n}\in\mathcal{Q}$ denote the negative features from a memory structure of size $K$.
+The loss associated with the query is computed using the logits $\ell(\mathbf{z_{i}})=\mathbf{q}^{T}\cdot\mathbf{z_{i}}/\tau$, which are processed through a softmax function.
+We define $\mathcal{\hat{Q}}=\{\mathbf{n}_{1},\mathbf{n}_{2},\ldots,\mathbf{n}_{K}\}$ as the ordered set of all negative features, where $\ell(\mathbf{n}_{i})>\ell(\mathbf{n}_{j})$ for all $i<j$, implying that the negative features are sorted based on decreasing similarity to the query.
+The most challenging negatives are selected by truncating the ordered set $\mathcal{\hat{Q}}$, retaining only the first $N<K$ elements, denoted as $\mathcal{\hat{Q}}^{N}$.
 Interpolated hard negatives (type 1).
-For each query $\mathbf{q}$, we propose to generate $N_{1}$ hard negative features by mixing the query $\mathbf{q}$ with a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$.
+For each query $\mathbf{q}$, we propose to generate $N_{1}$ synthetic hard negative features by mixing the query $\mathbf{q}$ with a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$.
 Let $H^{1}=\{\mathbf{h}_{1}^{1},\mathbf{h}_{2}^{1},\ldots,\mathbf{h}_{N_{1}}^{1}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic point $\mathbf{h}_{k}^{1}\in H^{1}$ would be given by:
+Then a synthetic negative feature $\mathbf{h}_{k}^{1}\in H^{1}$ would be given by:
 $\mathbf{h}_{k}^{1}=\alpha_{k}\cdot\mathbf{q}+(1-\alpha_{k})\cdot\mathbf{n}_{i},\quad\alpha_{k}\in(0,\alpha_{\text{max}})$ (4)
@@
 The resulting synthetic hard negatives are then normalized and added to the set of negative logits for the query.
-Interpolation creates a synthetic embedding that lies between the query and the negative in the embedding space.
-We set $\alpha_{\text{max}}=0.5$ to ensure that the query’s contribution is always smaller than the one of the negative.
-This is similar to the hardest negatives (type 2) of MoCHI.
+Interpolation creates a synthetic embedding that lies between the query and the negative in the representation space.
+We set $\alpha_{\text{max}}=0.5$ to guarantee that the contribution of the query is always less than that of the negative.
+This is similar to the hardest negatives (type 2) of MoCHI [23].
 Extrapolated hard negatives (type 2).
@@
 Let $H^{2}=\{\mathbf{h}_{1}^{2},\mathbf{h}_{2}^{2},\ldots,\mathbf{h}_{N_{2}}^{2}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic point $\mathbf{h}_{k}^{2}\in H^{2}$ would be given by:
+Then a synthetic negative feature $\mathbf{h}_{k}^{2}\in H^{2}$ would be given by:
 $\mathbf{h}_{k}^{2}=\mathbf{n}_{i}+\beta_{k}\cdot(\mathbf{n}_{i}-\mathbf{q}),\quad\beta_{k}\in(1,\beta_{\text{max}})$ (5)
@@
 Let $H^{3}=\{\mathbf{h}_{1}^{3},\mathbf{h}_{2}^{3},\ldots,\mathbf{h}_{N_{3}}^{3}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic point $\mathbf{h}_{k}^{3}\in H^{3}$ would be given by:
+Then a synthetic negative feature $\mathbf{h}_{k}^{3}\in H^{3}$ would be given by:
 $\mathbf{h}_{k}^{3}=\gamma_{k}\cdot\mathbf{n}_{i}+(1-\gamma_{k})\cdot\mathbf{n}_{j},\quad\gamma_{k}\in(0,1)$ (6)
@@
 Mixup combines pairs of the hardest existing negative features to create a synthetic embedding that represents a blend of challenging cases.
-This is similar to the hard negatives (type 1) of MoCHI.
+This is similar to the hard negatives (type 1) of MoCHI [23].
 Noise-injected hard negatives (type 4).
@@
 Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{4}=\{\mathbf{h}_{1}^{4},\mathbf{h}_{2}^{4},\ldots,\mathbf{h}_{N_{4}}^{4}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic point $\mathbf{h}_{k}^{4}\in H^{4}$ would be given by:
-$\mathbf{h}_{k}^{4}=\mathbf{n}_{i}+\mathcal{N}(0,\sigma^{2})$ (7)
-where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise with standard deviation $\sigma$.
+Then a synthetic negative feature $\mathbf{h}_{k}^{4}\in H^{4}$ would be given by:
+$\mathbf{h}_{k}^{4}=\mathbf{n}_{i}+\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ (7)
+where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ represents Gaussian noise with standard deviation $\sigma$ (where $\mathbf{I}$ is the identity matrix).
 The noisy negatives are normalized before being used in the loss calculation.
@@
 Given the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{5}=\{\mathbf{h}_{1}^{5},\mathbf{h}_{2}^{5},\ldots,\mathbf{h}_{N_{5}}^{5}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic point $\mathbf{h}_{k}^{5}\in H^{5}$ would be given by:
+Then a synthetic negative feature $\mathbf{h}_{k}^{5}\in H^{5}$ would be given by:
 $\mathbf{h}_{k}^{5}=\mathbf{n}_{i}+\delta\cdot\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i})$ (8)
 where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\text{sim}(\cdot,\cdot)$ is the similarity function and $\delta$ controls the perturbation magnitude.
-We approximate the gradients by calculating the difference in similarity scores with small perturbations applied.
 The perturbed embeddings are then normalized and added to the negative logits.
@@
 Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{6}=\{\mathbf{h}_{1}^{6},\mathbf{h}_{2}^{6},\ldots,\mathbf{h}_{N_{6}}^{6}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic point $\mathbf{h}_{k}^{6}\in H^{6}$ would be given by:
+Then a synthetic negative feature $\mathbf{h}_{k}^{6}\in H^{6}$ would be given by:
 $\mathbf{h}_{k}^{6}=\mathbf{n}_{i}+\eta\cdot\text{sign}(\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i}))$ (9)
@@
 Intuition of synthetic hard negatives.
-Each SynCo strategy enhances model generalization through challenging contrasts.
+Each SynCo strategy imrpoves model generalization through challenging contrasts.
 Type 1 interpolates between query and hard negatives, increasing sample diversity throughout training.
-Type 2 extrapolates beyond the query, pushing embedding space boundaries and improving robustness to difficult contrasts.
+Type 2 extrapolates beyond the query, pushing representation space boundaries and improving robustness to difficult contrasts.
 Type 3 combines pairs of hard negatives, encouraging more generalized and robust feature learning.
@@
 Type 6 applies adversarial perturbations, creating the most challenging contrasts to distinguish deceptively similar samples.
-A toy example of the proposed synthetic hard negative generation is presented in Figure 4.
-The latter figure visualizes the different types of synthetic hard negatives generated from a query point and three original negative points in a 2D embedding space.
-Dotted lines connect the query to extrapolated negatives and between pairs of mixup negatives, illustrating the relationships used to generate these synthetic points.
-Figure 4: A toy example illustrating the six types of synthetic hard negatives generated by SynCo from a query point and three original negative points in a 2D embedding space.
+A simplified example illustrating the synthetic hard negative generation is provided in Figure 4.
+The latter figure visualizes the different types of synthetic hard negatives generated from a query and three original negative features in a 2D representation space.
+Dotted lines connect the query to extrapolated negatives and between pairs of mixup negatives, illustrating the relationships used to generate these synthetic negative features.
+Figure 4: A toy example illustrating the six types of synthetic hard negatives generated by SynCo from a query and three original negative features in a 2D representation space.
 Is the proxy task more difficult?
-Figure 3 illustrates the proxy task performance for various configurations of SynCo.
+Figure 3 depicts the proxy task performance for different configurations of SynCo.
 We observe that incorporating synthetic negatives leads to faster learning and improved performance.
@@
 This indicates that SynCo presents the most challenging proxy task.
-The lower performance when including synthetic negatives suggests that $\max\ell(\mathbf{h}_{k}^{i})>\max\ell(\mathbf{n}_{j})$, $\mathbf{h}_{k}^{i}\in H^{i}$, $\mathbf{n}_{j}\in\tilde{\mathcal{Q}}_{N}$.
+This is evidenced by $\max\ell(\mathbf{h}_{k}^{i})>\max\ell(\mathbf{n}_{j})$, where $\mathbf{h}_{k}^{i}\in H^{i}$ are synthetic negatives and $\mathbf{n}_{j}\in\tilde{\mathcal{Q}}_{N}$ are original negatives.
 Through SynCo, we modulate proxy task difficulty via synthetic negatives, pushing the model to learn more robust features.
-Evaluating the usage of the embedding space.
-Wang and Isola [38] presented two losses/metrics for assessing contrastive learning representations.
-The first metric assesses alignment by measuring the absolute distance between representations that share the same label on a hypersphere.
-The second metric evaluates uniformity by calculating the logarithm of the average pairwise Gaussian potential among all embeddings.
-In Figure 5, we present these metrics for various models, based on re-implemented methods and experiments conducted using features from all images in the ImageNet-100 validation set.
-Our findings demonstrate that SynCo significantly enhances the uniformity of representations compared to MoCo-v2 and MoCHI, indicating its superior ability to utilize the embedding space in the proxy task.
+Evaluating the usage of the representation space.
+To assess learned representations, we employ alignment and uniformity metrics proposed by [44].
+These metrics provide insights into representation space utilization, with alignment quantifying the grouping of similar samples and uniformity measuring representation spread across the hypersphere.
+Figure 5 presents results for various models using features from the ImageNet-100 validation set.
+Our findings demonstrate that SynCo significantly enhances the uniformity of representations compared to MoCo-v2 and MoCHI, indicating its superior ability to utilize the representation space in the proxy task.
 Furthermore, the incorporation of synthetic negatives (types 1 to 6) leads to improved alignment.
@@
 Figure 5: Performance comparison of MoCo-v2, MoCHI, and SynCo (under various configurations) on ImageNet-100 in terms of alignment and uniformity metrics.
-The x/y axes correspond to $-\mathcal{L}_{\text{uniform}}$ and $-\mathcal{L}_{\text{align}}$, respectively.
-The best-performing model lies towards the top-right of the plot.
+The x-axis and y-axis represent $-\mathcal{L}_{\text{uniform}}$ and $-\mathcal{L}_{\text{align}}$, respectively.
+The model with the highest performance is located in the upper-right corner of the chart.
 We use $K=65$k.
@@
 5.1 Implementation Details
-We learn representations on the common ImageNet ILSVRC-2012 [9] and its smaller ImageNet-100 subset [32] using a ResNet-50 [15] encoder.
-All runs of SynCo are based on MoCo-v2 [8].
-We developed our approach on top of the official public implementation of MoCo and reproduced it on our setup.
-We run all experiments on 4 NVIDIA A40 GPUs.
-For linear classification on ImageNet, we follow the common protocol and report results on the validation set.
-We report performance after learning linear classifiers for 100 epochs, with an initial learning rate of $30.0$ ($10.0$ for ImageNet-100), a batch size of 256 and a cosine learning rate schedule.
+We pretrain SynCo on ImageNet ILSVRC-2012 [10] and its smaller ImageNet-100 subset [38] for ablation studies using a ResNet-50 [18] encoder.
+Our implementation builds upon MoCo-v2 [8].
 For training, unless stated otherwise, we use $K=65$k ($K=16$k for ImageNet-100).
@@
 For hard negative generation, we use the top $N=1024$ hardest negatives, with $N_{1}=N_{2}=N_{3}=256$ and $N_{4}=N_{5}=N_{6}=64$.
-For ImageNet, we report accuracy for a single-crop testing.
-For object detection on PASCAL VOC [11] we follow [17] and fine-tune a Faster R-CNN [30], R50-C4 on trainval07+12 and test on test2007.
-For COCO object detection, we follow [17] and fine-tune a Mask R-CNN [16], R50-C4 on train2017 and test on val2017.
-We use the open-source Detectron2 [40] code and report the common $AP$, $AP_{50}$ and $AP_{75}$ metrics.
-Similar to [17], we do not perform hyperparameter tuning for the object detection task.
+For ImageNet linear evaluation, we train a linear classifier on frozen features for 100 epochs, using a batch size of 256 and a cosine learning rate schedule.
+Initial learning rates are set to 30.0 for ImageNet and 10.0 for ImageNet-100.
+To evaluate transfer learning, we apply SynCo to object detection tasks.
+For PASCAL VOC [12], we fine-tune a Faster R-CNN [35] (R50-C4) on trainval07+12 and test on test2007.
+For COCO [28], we use a Mask R-CNN [19] (R50-C4), fine-tuning on train2017 and evaluating on val2017.
+We employ Detectron2 [46] and report standard AP metrics, following [20] without additional hyperparameter tuning.
 Detailed implementation details are provided in the supplementary material.
 5.2 Linear Evaluation on ImageNet
-We evaluate the SynCo representation by training a linear classifier on top of the frozen features that were pretrained on ImageNet, following the procedure described in [23, 22, 6, 13, 36].
-Full details of the linear evaluation protocol are described in the supplementary material.
-With a standard ResNet-50 and for 200 epoch pretraining, SynCo obtains $67.9\%\pm 0.16\%$ top-1 accuracy and $88.0\%\pm 0.05\%$ top-5 accuracy (Table 1).
-Specifically, SynCo achieves a +0.4% top-1 accuracy improvement over MoCo-v2 and a +1.0% improvement over MoCHI with 200 epochs of pretraining, leveraging synthetic hard negatives.
+We evaluate the SynCo representation by training a linear classifier on top of the frozen features that were pretrained on ImageNet, following the procedure described in [26, 25, 6, 14, 42] (details in supplementary material).
+SynCo obtains $67.9\%\pm 0.16\%$ top-1 accuracy and $88.0\%\pm 0.05\%$ top-5 accuracy after 200 epoch pretraining (Table 1).
+Specifically, SynCo achieves a +0.4% top-1 accuracy improvement over MoCo-v2 and a +1.0% improvement over MoCHI.
 While MoCHI, which employs hard negatives, achieves lower performance than MoCo-v2, our method, which generates synthetic hard negatives, not only avoids this drop in performance but actually improves it.
@@
 When training for 800 epochs, SynCo obtains 70.6% top-1 accuracy (89.8% top-5 accuracy) (Table 2).
-This represents a +1.9% improvement in top-1 accuracy over MoCHI and a -0.5% decrease compared to MoCo-v2 under the same 800 epochs pretraining settings.
+This represents a +1.9% improvement in top-1 accuracy over MoCHI.
 These results show that SynCo achieves higher accuracy than its competitors with only a minor computational overhead for generating synthetic hard negatives, while also enabling faster training convergence.
@@
 5.3 Semi-supervised Training on ImageNet
-We evaluate the performance obtained when fine-tuning SynCo representation, pretrained for 800 epochs, on a classification task with a small subset of ImageNet’s train set, this time using label information.
-We follow the semi-supervised protocol of [23, 45, 6, 13, 36], and use the same fixed splits of respectively 1% and 10% of ImageNet labeled training data as in [6].
-Our complete semi-supervised learning setup is detailed in the supplementary material.
-Despite not being specifically designed for this task, SynCo shows promising results (Table 3).
-With 1% labeled data, SynCo achieves 50.8% $\pm$ 0.21% top-1 accuracy and 77.5% $\pm$ 0.12% top-5 accuracy.
-With 10% labeled data, SynCo reaches 66.6% $\pm$ 0.19% top-1 accuracy and 88.0% $\pm$ 0.10% top-5 accuracy.
+We evaluate SynCo in a semi-supervised setting using 1% and 10% of labeled ImageNet data, following protocols from [26, 52, 6, 14, 42] (details in supplementary material).
+We use the same data splits as [6].
+Results in Table 3 show SynCo’s competitive performance.
+With 1% labeled data, SynCo achieves $50.8\%\pm 0.21\%$ top-1 accuracy and $77.5\%\pm 0.12\%$ top-5 accuracy.
+Using 10% labeled data, it reaches $66.6\%\pm 0.19\%$ top-1 accuracy and $88.0\%\pm 0.10\%$ top-5 accuracy.
 We observe that despite a lower volume of labeled data, SynCo competes well with established semi-supervised methods like SimCLR, BYOL, and Barlow Twins.
@@
 We evaluate the SynCo representation, pretrained for 200 epochs, by applying it to the detection task.
-We follow the protocol of [17], and as used in [20] for a fair comparison.
+We follow the protocol of [20].
 Detailed configurations for object detection experiments are provided in in the supplementary material.
@@
 MoCHI 57.5 82.7 64.4 39.2 58.9 42.4 34.4 55.5 36.6
-SynCo (ours) 57.2 82.5 63.9 40.4 60.1 43.8 35.4 57.0 37.4
+SynCo (ours) 57.2 82.6 63.9 41.0 60.6 44.8 35.7 57.4 38.1
 5.5 Ablation Study
@@
 6 Conclusion
-In this paper, we built upon the observation that harder negatives are needed for more effective self-supervised contrastive learning.
-Based on that, we presented SynCo, a synthetic hard negative generation approach that improves the quality of representations learned in an unsupervised way, offering better transfer learning performance and utilization of the embedding space.
-We showed that our method learns generalizable representations, which is important considering the high compute cost of self-supervised learning.
-We found that multiple SynCo configurations provide considerable gains, and that using synthetic hard negatives consistently improves transfer learning performance.
+This paper introduces SynCo, a novel approach to contrastive learning that leverages synthetic hard negatives to enhance visual representation learning.
+By generating diverse and challenging negative samples on-the-fly, SynCo addresses the limitations of existing methods in maintaining a pool of effective hard negatives throughout training.
+Our comprehensive experiments demonstrate that SynCo not only accelerates the learning process but also leads to more robust and transferable representations.
+The effectiveness of SynCo is evidenced by its superior performance across various benchmarks, including linear evaluation on ImageNet, semi-supervised learning tasks, and transfer learning to object detection on PASCAL VOC and COCO datasets.
 While our experiments primarily utilized the MoCo framework, the proposed hard negative generation strategies are general and can be applied to any contrastive learning method that benefits from hard negatives, including SimCLR[6], CPC [42], PIRL [30], and others [49, 45, 24, 11, 40].
 These methods, which employ the InfoNCE loss function (or its variants [6, 11]) and instance discrimination as the pretext task, can benefit from the enhanced hard negative generation strategies proposed by SynCo.
-The introduction of synthetic hard negatives can provide these methods with more challenging and informative contrasts, potentially leading to better feature representations.
-Also, the applicability of SynCo extends beyond visual representation learning.
+The introduction of synthetic hard negatives can provide these methods with more challenging, informative contrasts, potentially leading to better feature representations.
+Also, SynCo’s applicability extends beyond visual representation learning.
 The concept of hard negatives can be applied to various domains such as natural language processing, audio processing, and other areas where contrastive learning is relevant.
+Broader Impact
+The presented research should be categorized as research in the field of unsupervised learning.
+This work may inspire new algorithms, theoretical, and experimental investigation.
+The algorithm presented here can be used for many different vision applications and a particular use may have both positive or negative impacts, which is known as the dual use problem.
+Besides, as vision datasets could be biased, the representation learned by SynCo could be susceptible to replicate these biases.

### v4

@@
 Building on the MoCo framework, SynCo introduces six strategies for creating diverse synthetic hard negatives on-the-fly with minimal computational overhead.
-SynCo achieves faster training and better representation learning, reaching 67.9% top-1 accuracy on ImageNet ILSVRC-201 linear evaluation after 200 pretraining epochs, surpassing MoCo’s 67.5% using the same ResNet-50 encoder.
-It also transfers more effectively to detection tasks: on PASCAL VOC, it outperforms both the supervised baseline and MoCo with 82.6% AP; on COCO, it sets new benchmarks with 41.0% AP for bounding box detection and 35.7% AP for instance segmentation.
+SynCo achieves faster training and better representation learning, reaching 67.9% top-1 accuracy on ImageNet ILSVRC-2012 linear evaluation after 200 pretraining epochs, surpassing MoCo’s 67.5% using the same ResNet-50 encoder.
+It also transfers more effectively to detection tasks: on PASCAL VOC, it outperforms both the supervised baseline and MoCo with 82.5% AP; on COCO, it sets new benchmarks with 40.9% AP for bounding box detection and 35.5% AP for instance segmentation.
 Our synthetic hard negative generation approach significantly enhances visual representations learned through self-supervised contrastive learning.
@@
 1 Introduction
-Figure 1: SynCo extends MoCo by introducing synthetic hard negatives generated on-the-fly from a memory queue.
+Figure 1: SynCo extends MoCo [21, 9] by introducing synthetic hard negatives generated on-the-fly from a memory queue.
 The process begins with two augmented views of an image, $\mathbf{x}_{q}$ and $\mathbf{x}_{k}$, processed by an encoder and a momentum encoder, respectively, producing feature vectors $\mathbf{q}$ and $\mathbf{k}$.
-The memory queue holds negative samples $\mathbf{n}_{1},\mathbf{n}_{2},\ldots$, which are concatenated with synthetic hard negatives $\mathbf{h}_{1},\mathbf{h}_{2},\ldots$ generated using the SynCo strategies.
+The memory queue holds negative samples $\mathbf{n}_{1},\mathbf{n}_{2},\ldots$, which are concatenated with synthetic hard negatives $\mathbf{s}_{1},\mathbf{s}_{2},\ldots$ generated using the SynCo strategies.
 These combined negatives are used to compute the affinity matrix, which, together with the positive pair (query $\mathbf{q}$ and key $\mathbf{k}$), contributes to the InfoNCE loss calculation.
@@
   2. 2.
-    We empirically show improved downstream performance on ImageNet by incorporating synthetic hard negatives, demonstrating enhancements in both linear evaluation and semi-supervised learning tasks.
+    We empirically show improved downstream performance on ImageNet ILSVRC-2012 by incorporating synthetic hard negatives, demonstrating improvements in both linear evaluation and semi-supervised learning tasks.
   3. 3.
-    We show SynCo learns stronger representations by measuring their transfer learning capabilities COCO and PASCAL VOC detection, where it outperforms both the supervised baseline and MoCo.
+    We show that SynCo learns stronger representations by measuring their transfer learning capabilities COCO and PASCAL VOC detection, where it outperforms both the supervised baseline and MoCo.
 The paper is structured as follows: Section 2 reviews related work; Section 3 explores hard negatives in contrastive learning; Section 4 introduces our synthetic hard negatives method; Section 5 presents experimental results; and Section 6 concludes the paper.
@@
 Building on these ideas, MoCHI [24] has explored integrating hard negative mixing into existing frameworks to further improve performance.
-When incorporating these approaches into representation space generation, it’s essential to consider how negative samples impact the model’s learning space.
-The goal is to implement a strategy that not only challenges the model but also enhances its ability to extract and differentiate features effectively.
 By employing these methods, models become more adept at handling detailed and complex tasks, ensuring each negative sample significantly contributes to optimizing learning outcomes and boosting overall model effectiveness.
@@
 By adding synthetic features to the representation space, models improve in generalization and performance.
-Methods have generated features for unseen classes using generative models [17, 48, 36], while others have integrated these into self-supervised and contrastive learning frameworks [27, 54].
+Some methods have generated features for unseen classes using generative models [18, 49, 37], while others have integrated these into self-supervised and contrastive learning frameworks [28, 55].
 Models that combine generation with representation spaces have also shown success in zero-shot learning [17].
@@
 4 Synthetic Hard Negatives in Contrastive Learning
-Based on these observations, we present an approach for generating synthetic hard negatives in the representation space using six distinct strategies.
-We refer to the proposed approach as SynCo (”Synthetic negatives in Contrastive learning”).
-4.1 Hard Negatives Generation
+In this section, we present an approach for generating synthetic hard negatives in the representation space using six distinct strategies.
+We refer to our proposed approach as SynCo (”Synthetic negatives in Contrastive learning”).
+4.1 Generating Synthetic Hard Negatives
 Let $\mathbf{q}$ represent the query image, $\mathbf{k}$ its corresponding key, and $\mathbf{n}\in\mathcal{Q}$ denote the negative features from a memory structure of size $K$.
@@
 The most challenging negatives are selected by truncating the ordered set $\mathcal{\hat{Q}}$, retaining only the first $N<K$ elements, denoted as $\mathcal{\hat{Q}}^{N}$.
-Interpolated hard negatives (type 1).
+Interpolated synthetic negatives (type 1).
 For each query $\mathbf{q}$, we propose to generate $N_{1}$ synthetic hard negative features by mixing the query $\mathbf{q}$ with a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$.
-Let $H^{1}=\{\mathbf{h}_{1}^{1},\mathbf{h}_{2}^{1},\ldots,\mathbf{h}_{N_{1}}^{1}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic negative feature $\mathbf{h}_{k}^{1}\in H^{1}$ would be given by:
-$\mathbf{h}_{k}^{1}=\alpha_{k}\cdot\mathbf{q}+(1-\alpha_{k})\cdot\mathbf{n}_{i},\quad\alpha_{k}\in(0,\alpha_{\text{max}})$ (4)
+Let $S^{1}=\{\mathbf{s}_{1}^{1},\mathbf{s}_{2}^{1},\ldots,\mathbf{s}_{N_{1}}^{1}\}$ be the set of synthetic negatives to be generated.
+Then a synthetic negative feature $\mathbf{s}_{k}^{1}\in S^{1}$ would be given by:
+$\mathbf{s}_{k}^{1}=\alpha_{k}\cdot\mathbf{q}+(1-\alpha_{k})\cdot\mathbf{n}_{i},\quad\alpha_{k}\in(0,\alpha_{\text{max}})$ (4)
 where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\alpha_{k}$ is randomly sampled from a uniform distribution in the range $(0,\alpha_{\text{max}})$.
@@
 This is similar to the hardest negatives (type 2) of MoCHI [24].
-Extrapolated hard negatives (type 2).
+Extrapolated synthetic negatives (type 2).
 For each query $\mathbf{q}$, we propose to generate $N_{2}$ hard negative features by extrapolating beyond the query embedding in the direction of the hardest negative features.
 Similar to the interpolated method, we use a randomly chosen feature from the $N$ hardest negatives in $\mathcal{\hat{Q}}^{N}$.
-Let $H^{2}=\{\mathbf{h}_{1}^{2},\mathbf{h}_{2}^{2},\ldots,\mathbf{h}_{N_{2}}^{2}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic negative feature $\mathbf{h}_{k}^{2}\in H^{2}$ would be given by:
-$\mathbf{h}_{k}^{2}=\mathbf{n}_{i}+\beta_{k}\cdot(\mathbf{n}_{i}-\mathbf{q}),\quad\beta_{k}\in(1,\beta_{\text{max}})$ (5)
+Let $S^{2}=\{\mathbf{s}_{1}^{2},\mathbf{s}_{2}^{2},\ldots,\mathbf{s}_{N_{2}}^{2}\}$ be the set of synthetic negatives to be generated.
+Then a synthetic negative feature $\mathbf{s}_{k}^{2}\in S^{2}$ would be given by:
+$\mathbf{s}_{k}^{2}=\mathbf{n}_{i}+\beta_{k}\cdot(\mathbf{n}_{i}-\mathbf{q}),\quad\beta_{k}\in(1,\beta_{\text{max}})$ (5)
 where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\beta_{k}$ is randomly sampled from a uniform distribution in the range $(1,\beta_{\text{max}})$.
@@
 We choose $\beta_{\text{max}}=1.5$.
-Mixup hard negatives (type 3).
+Mixup synthetic negatives (type 3).
 For each query $\mathbf{q}$, we propose to generate $N_{3}$ hard negative features by combining pairs of the $N$ hardest existing negative features in $\mathcal{\hat{Q}}^{N}$.
-Let $H^{3}=\{\mathbf{h}_{1}^{3},\mathbf{h}_{2}^{3},\ldots,\mathbf{h}_{N_{3}}^{3}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic negative feature $\mathbf{h}_{k}^{3}\in H^{3}$ would be given by:
-$\mathbf{h}_{k}^{3}=\gamma_{k}\cdot\mathbf{n}_{i}+(1-\gamma_{k})\cdot\mathbf{n}_{j},\quad\gamma_{k}\in(0,1)$ (6)
+Let $S^{3}=\{\mathbf{s}_{1}^{3},\mathbf{s}_{2}^{3},\ldots,\mathbf{s}_{N_{3}}^{3}\}$ be the set of synthetic negatives to be generated.
+Then a synthetic negative feature $\mathbf{s}_{k}^{3}\in S^{3}$ would be given by:
+$\mathbf{s}_{k}^{3}=\gamma_{k}\cdot\mathbf{n}_{i}+(1-\gamma_{k})\cdot\mathbf{n}_{j},\quad\gamma_{k}\in(0,1)$ (6)
 where $\mathbf{n}_{i},\mathbf{n}_{j}\in\mathcal{\hat{Q}}^{N}$ and $\gamma_{k}$ is randomly sampled from a uniform distribution in the range $(0,1)$.
@@
 This is similar to the hard negatives (type 1) of MoCHI [24].
-Noise-injected hard negatives (type 4).
+Noise-injected synthetic negatives (type 4).
 For each query $\mathbf{q}$, we propose to generate $N_{4}$ hard negative features by adding Gaussian noise to the hardest negative features.
-Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{4}=\{\mathbf{h}_{1}^{4},\mathbf{h}_{2}^{4},\ldots,\mathbf{h}_{N_{4}}^{4}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic negative feature $\mathbf{h}_{k}^{4}\in H^{4}$ would be given by:
-$\mathbf{h}_{k}^{4}=\mathbf{n}_{i}+\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ (7)
+Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $S^{4}=\{\mathbf{s}_{1}^{4},\mathbf{s}_{2}^{4},\ldots,\mathbf{s}_{N_{4}}^{4}\}$ be the set of synthetic negatives to be generated.
+Then a synthetic negative feature $\mathbf{s}_{k}^{4}\in S^{4}$ would be given by:
+$\mathbf{s}_{k}^{4}=\mathbf{n}_{i}+\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ (7)
 where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\mathcal{N}(\mathbf{0},\sigma^{2}\cdot\mathbf{I})$ represents Gaussian noise with standard deviation $\sigma$ (where $\mathbf{I}$ is the identity matrix).
@@
 Noise injection adds Gaussian noise to the hardest negative features, resulting in a synthetic embedding with added randomness.
-Perturbed hard negatives (type 5).
+Perturbed synthetic negatives (type 5).
 For each query $\mathbf{q}$, we propose to generate $N_{5}$ hard negative features by perturbing the embeddings of the hardest negative features.
-Given the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{5}=\{\mathbf{h}_{1}^{5},\mathbf{h}_{2}^{5},\ldots,\mathbf{h}_{N_{5}}^{5}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic negative feature $\mathbf{h}_{k}^{5}\in H^{5}$ would be given by:
-$\mathbf{h}_{k}^{5}=\mathbf{n}_{i}+\delta\cdot\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i})$ (8)
+Given the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $S^{5}=\{\mathbf{s}_{1}^{5},\mathbf{s}_{2}^{5},\ldots,\mathbf{s}_{N_{5}}^{5}\}$ be the set of synthetic negatives to be generated.
+Then a synthetic negative feature $\mathbf{s}_{k}^{5}\in S^{5}$ would be given by:
+$\mathbf{s}_{k}^{5}=\mathbf{n}_{i}+\delta\cdot\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i})$ (8)
 where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\text{sim}(\cdot,\cdot)$ is the similarity function and $\delta$ controls the perturbation magnitude.
@@
 Perturbation modifies the embeddings of the hardest negative features based on the gradient of the similarity function, creating synthetic negatives that are slightly adjusted to be more challenging for the model.
-Adversarial hard negatives (type 6).
+Adversarial synthetic negatives (type 6).
 For each query $\mathbf{q}$, we propose to generate $N_{6}$ hard negative features by applying adversarial perturbations to the hardest negative features to maximize their similarity to the query embeddings.
-Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $H^{6}=\{\mathbf{h}_{1}^{6},\mathbf{h}_{2}^{6},\ldots,\mathbf{h}_{N_{6}}^{6}\}$ be the set of synthetic negatives to be generated.
-Then a synthetic negative feature $\mathbf{h}_{k}^{6}\in H^{6}$ would be given by:
-$\mathbf{h}_{k}^{6}=\mathbf{n}_{i}+\eta\cdot\text{sign}(\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i}))$ (9)
+Using the top $N$ hardest negatives $\mathcal{\hat{Q}}^{N}$, let $S^{6}=\{\mathbf{s}_{1}^{6},\mathbf{s}_{2}^{6},\ldots,\mathbf{s}_{N_{6}}^{6}\}$ be the set of synthetic negatives to be generated.
+Then a synthetic negative feature $\mathbf{s}_{k}^{6}\in S^{6}$ would be given by:
+$\mathbf{s}_{k}^{6}=\mathbf{n}_{i}+\eta\cdot\text{sign}(\nabla_{\mathbf{n}_{i}}\text{sim}(\mathbf{q},\mathbf{n}_{i}))$ (9)
 where $\mathbf{n}_{i}\in\mathcal{\hat{Q}}^{N}$ and $\eta$ controls the perturbation magnitude.
@@
 Adversarial hard negatives apply adversarial perturbations to the hardest negative features, specifically altering them to maximize their similarity to the query embeddings, thereby producing the most challenging contrasts.
+4.2 Integrating Synthetic Hard Negatives into the Contrastive Loss
+The synthetic hard negatives generated are integrated into the contrastive learning process by modifying the InfoNCE loss.
+Let $\mathcal{S}=\bigcup_{i=1}^{6}S^{i}$ represent the concatenation of all synthetic hard negatives, where $S^{i}$ is the set of synthetic negatives generated by the $i$-th strategy.
+This combined set of synthetic negatives augments the original negatives $\mathcal{Q}$, providing a more diverse and challenging set of contrasts for the query.
+The modified InfoNCE loss is given by:
+$\mathcal{L}(\mathbf{q},\mathbf{k},\mathcal{Q},\mathcal{S})=-\log\frac{\exp(\mathbf{q}^{T}\cdot\mathbf{k}/\tau)}{Z+\sum_{\mathbf{s}\in\mathcal{S}}\exp(\mathbf{q}^{T}\cdot\mathbf{s}/\tau)}$ (10)
+where $Z=\exp(\mathbf{q}^{T}\cdot\mathbf{k}/\tau)+\sum_{\mathbf{n}\in\mathcal{Q}}\exp(\mathbf{q}^{T}\cdot\mathbf{n}/\tau)$, $\tau$ is the temperature parameter, $\mathcal{Q}$ is the set of original negatives, and $\mathcal{S}$ is the set of synthetic hard negatives.
+By incorporating both real and synthetic negatives, the model is exposed to a wider variety of challenging examples, which encourages learning more robust and generalizable representations.
+The overall computational overhead of SynCo is roughly equivalent to increasing the queue/memory by $\sum_{i=1}^{6}N_{i}\ll K$, along with the additional cost of generating the synthetic negatives.
 4.3 Discussion on Synthetic Hard Negatives
-Intuition of synthetic hard negatives.
-Each SynCo strategy imrpoves model generalization through challenging contrasts.
-Type 1 interpolates between query and hard negatives, increasing sample diversity throughout training.
-Type 2 extrapolates beyond the query, pushing representation space boundaries and improving robustness to difficult contrasts.
-Type 3 combines pairs of hard negatives, encouraging more generalized and robust feature learning.
-Type 4 injects Gaussian noise, promoting invariance to minor feature fluctuations and enhancing generalization.
-Type 5 modifies embeddings based on similarity gradients, refining discriminatory power by directing the model towards harder negatives.
-Type 6 applies adversarial perturbations, creating the most challenging contrasts to distinguish deceptively similar samples.
-A simplified example illustrating the synthetic hard negative generation is provided in Figure 4.
-The latter figure visualizes the different types of synthetic hard negatives generated from a query and three original negative features in a 2D representation space.
-Dotted lines connect the query to extrapolated negatives and between pairs of mixup negatives, illustrating the relationships used to generate these synthetic negative features.
-Figure 4: A toy example illustrating the six types of synthetic hard negatives generated by SynCo from a query and three original negative features in a 2D representation space.
+The following explores the effects of incorporating synthetic hard negatives on the difficulty of the proxy task and how they influence the usage of the representation space.
 Is the proxy task more difficult?
@@
 This indicates that SynCo presents the most challenging proxy task.
-This is evidenced by $\max\ell(\mathbf{h}_{k}^{i})>\max\ell(\mathbf{n}_{j})$, where $\mathbf{h}_{k}^{i}\in H^{i}$ are synthetic negatives and $\mathbf{n}_{j}\in\tilde{\mathcal{Q}}_{N}$ are original negatives.
+This is evidenced by $\max\ell(\mathbf{s}_{k}^{i})>\max\ell(\mathbf{n}_{j})$, where $\mathbf{s}_{k}^{i}\in S^{i}$ are synthetic negatives and $\mathbf{n}_{j}\in\tilde{\mathcal{Q}}_{N}$ are original negatives.
 Through SynCo, we modulate proxy task difficulty via synthetic negatives, pushing the model to learn more robust features.
@@
 To evaluate transfer learning, we apply SynCo to object detection tasks.
-For PASCAL VOC [12], we fine-tune a Faster R-CNN [35] (R50-C4) on trainval07+12 and test on test2007.
-For COCO [28], we use a Mask R-CNN [19] (R50-C4), fine-tuning on train2017 and evaluating on val2017.
+For PASCAL VOC [13], we fine-tune a Faster R-CNN [36] on trainval07+12 and test on test2007.
+For COCO [29], we use a Mask R-CNN [20], fine-tuning on train2017 and evaluating on val2017.
 We employ Detectron2 [47] and report standard AP metrics, following [21] without additional hyperparameter tuning.
@@
 Supervised 76.5 -
-PIRL 63.6 -
-LA 60.2 -
-InfoMin 70.1 89.4
-SimSiam 68.1 -
-SimCLR 69.3 -
-MoCo 60.7 -
-MoCo-v2 67.5 90.1
-MoCHI 66.9 -
+PIRL [31] 63.6 -
+LA [56] 60.2 -
+InfoMin [40] 70.1 89.4
+SimSiam [8] 68.1 -
+SimCLR [7] 69.3 -
+MoCo [21] 60.7 -
+MoCo-v2 [9] 67.5 90.1
+MoCHI [24] 66.9 -
 SynCo (ours) 67.9 88.0
@@
 Method Epochs Top-1 Top-5
-PIRL 800 63.6 -
-InfoMin 800 73.0 91.1
-BYOL 1000 74.3 91.6
-SimSiam 800 68.1 -
-SimCLR 1000 69.3 -
-Barlow Twins 1000 73.2 91.0
-MoCo-v2 800 71.1 90.1
-MoCHI 800 68.7 -
+PIRL [31] 800 63.6 -
+InfoMin [40] 800 73.0 91.1
+BYOL [15] 1000 74.3 91.6
+SimSiam [8] 800 68.1 -
+SimCLR [7] 1000 69.3 -
+BT [52] 1000 73.2 91.0
+MoCo-v2 [9] 800 71.1 90.1
+MoCHI [24] 800 68.7 -
 SynCo (ours) 800 70.6 89.8
@@
 Supervised 25.4 56.4 48.4 80.4
-InstDis 200 - - 39.2 77.4
-PIRL 800 30.7 60.4 57.2 83.8
-SimCLR 1000 48.3 65.6 75.5 87.8
-Barlow Twins 1000 55.0 69.7 79.2 89.3
-BYOL 1000 53.2 68.8 78.4 89.0
-SwAV 800 53.9 70.2 78.5 89.9
+InstDis [48] 200 - - 39.2 77.4
+PIRL [31] 800 30.7 60.4 57.2 83.8
+SimCLR [7] 1000 48.3 65.6 75.5 87.8
+BT [52] 1000 55.0 69.7 79.2 89.3
+BYOL [15] 1000 53.2 68.8 78.4 89.0
+SwAV [5] 800 53.9 70.2 78.5 89.9
 SynCo (ours) 800 50.8 66.6 77.5 88.0
@@
 Supervised 53.5 81.3 58.8 38.2 58.2 41.6 33.3 54.7 35.2
-InstDis - - - 37.4 57.6 40.6 34.1 54.6 36.4
-PIRL 55.5 81.0 61.3 38.5 57.6 41.2 34.0 54.6 36.2
-InfoMin 57.6 82.7 64.6 39.0 58.5 42.0 34.1 55.2 36.3
-SwAV 56.1 82.6 62.7 38.2 58.2 41.3 33.8 55.2 35.9
-SimSiam 57.0 82.4 63.7 39.2 59.3 42.1 34.4 56.0 36.7
-Barlow Twins 56.8 82.6 63.4 39.2 59.0 42.5 34.3 56.0 36.5
-MoCo 55.5 81.5 61.3 38.5 58.3 41.6 33.6 54.8 35.6
-MoCo-v2 57.0 82.4 63.6 39.0 58.6 41.9 34.2 55.4 36.2
-MoCHI 57.5 82.7 64.4 39.2 58.9 42.4 34.4 55.5 36.6
+InstDis [48] - - - 37.4 57.6 40.6 34.1 54.6 36.4
+PIRL [31] 55.5 81.0 61.3 38.5 57.6 41.2 34.0 54.6 36.2
+InfoMin [40] 57.6 82.7 64.6 39.0 58.5 42.0 34.1 55.2 36.3
+SwAV [5] 56.1 82.6 62.7 38.2 58.2 41.3 33.8 55.2 35.9
+SimSiam [8] 57.0 82.4 63.7 39.2 59.3 42.1 34.4 56.0 36.7
+BT [52] 56.8 82.6 63.4 39.2 59.0 42.5 34.3 56.0 36.5
+MoCo [21] 55.5 81.5 61.3 38.5 58.3 41.6 33.6 54.8 35.6
+MoCo-v2 [9] 57.0 82.4 63.6 39.0 58.6 41.9 34.2 55.4 36.2
+MoCHI [24] 57.5 82.7 64.4 39.2 58.9 42.4 34.4 55.5 36.6
 SynCo (ours) 57.2 82.6 63.9 41.0 60.6 44.8 35.7 57.4 38.1
@@
 Additional ablation studies and analyses are presented in the supplementary material.
-Type of hard negative.
+Ablation on type of hard negative.
 We evaluate the impact of each synthetic hard negative type on pretraining.
@@
 We see that every SynCo configuration outperform the MoCo-v2 baseline.
-Parameter sensitivity analysis.
+Ablation on hyperparameters.
 We conducted ablations on the parameters $\sigma$, $\delta$, and $\eta$ of SynCo’s type 4, type 5, and type 6 negatives, respectively.
@@
 This suggests that SynCo is robust across a wide range of values for $\sigma$, $\delta$, $\eta$.
-Queue size.
+Ablation on queue size.
 We investigate the effect of queue size $\mathcal{Q}$ on performance.
@@
 The concept of hard negatives can be applied to various domains such as natural language processing, audio processing, and other areas where contrastive learning is relevant.
-Broader Impact
-The presented research should be categorized as research in the field of unsupervised learning.
-This work may inspire new algorithms, theoretical, and experimental investigation.
-The algorithm presented here can be used for many different vision applications and a particular use may have both positive or negative impacts, which is known as the dual use problem.
-Besides, as vision datasets could be biased, the representation learned by SynCo could be susceptible to replicate these biases.
+Acknowledgments
+We would like to express our gratitude to Andreas Floros for his valuable feedback, particularly his assistance with equations, notations, and insightful discussions that greatly contributed to this work.